# Parameter inference of microlensed gravitational waves using neural spline flows

### v1

@@ Abstract
When gravitational waves (GWs) propagate near massive objects, they undergo gravitational lensing that imprints lens model dependent modulations on the waveform. This effect provides a powerful tool for cosmological and astrophysical studies. However, conventional Bayesian parameter inference methods for GWs are computationally expensive, especially for lensed events with additional lens parameters, necessitating more efficient approaches. In this work, we explore the use of neural spline flows (NSFs) for posterior inference of microlensed GWs, and successfully apply NSFs to the inference of 13-dimensional lens parameters. Our results demonstrate that compared with traditional methods like Bilby dynesty that rely on Bayesian inference, the NSF network we built not only achieves inference accuracy comparable to traditional methods for the main parameters ($M_{\rm L}$, $D_{\rm L}$, $Z_{\rm L}$, $\mathcal{M}_{\rm c}$, $d_{\rm L}$, $Z_{\rm S}$, $\iota$), but also can reduce the inference time from approximately 3 days to 0.8 s on average. Additionally, the network exhibits strong generalization for the spin parameters of GW sources. It is anticipated to become a powerful tool for future low-latency searches for lensed GW signals.
@@ I Introduction
Since the first direct detection of gravitational waves (GWs) in 2015, LIGO-Virgo-KAGRA (LVK) network have reported $93$ GW events in its first three observing runs, which enables us to conduct various novel tests of general relativity Abbott et al. (2016, 2018, 2019a, 2019b, 2021a) and cosmological researches The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2023). With the further analysis of the data observed in the fourth observing run of the LVK network and the construction of various high-performance gravitational wave detectors in the future (such as the Laser Interferometer Space Antenna (LISA) Barausse et al. (2020), Taiji Ruan et al. (2020), TianQin Luo et al. (2025), Einstein Telescope (ET) Maggiore et al. (2020), and Cosmic Explorer (CE) Evans et al. (2021)), the number of detected GW events will increase sharply, providing a brand-new powerful tool for research in fields such as fundamental physics, cosmology, and astrophysics. Among these GW signals, it is particularly worthy of attention those signals affected by the lensing effect. The lensing effect of GWs is caused by celestial bodies located between detectors and sources, which will result in a significant modulation of GW signals Dai and Venumadhav ; Dai et al. (2018).
@@
Recently, multiple studies have conducted numerous searches for the GW lensing effect, but none have clearly detected lensed GW signals. The methods employed include traditional search techniques based on matched filtering Abbott et al. (2021b, 2024); Janquart et al. (2023); Barsode et al. (2025); Goyal et al. (2024) and those incorporating deep learning technology Goyal et al. (2021); Kim et al. (2022); Schäfer et al. (2023). Nevertheless, as more and more GW events are expected to be reported, and given the predictions regarding the detection rate Li et al. (2018); Ng et al. (2018); Wang et al. (2021b); Xu et al. (2022); Brando et al. (2025), we remain hopeful that GW signals affected by the gravitational lensing effect will be detectable in future observational runs.
@@
Parameter inference for lensed GW events is a crucial step in studying the GW lensing effect. The traditional parameter inference process based on Bayesian methods is extremely time-consuming, especially for lensed GW events with additional lens parameters, which significantly reduces its efficiency. Recent studies have shown that deep learning techniques such as Conditional Variational Autoencoders (CVAE) Gabbard et al. (2022) and Normalizing Flows (NFs) Green and Gair (2021); Green et al. (2020); Dax et al. (2021); Shen et al. (2022); Whittaker et al. (2022); Dax et al. (2023); Zhao et al. (2023); Sun et al. (2024a); Du et al. (2024); Sun et al. (2024b); Leyde et al. (2024); Dax et al. (2025); Xiong et al. (2025) can effectively and rapidly perform parameter inference, providing powerful tools for low-latency GW signal searches. In Ref. Nerin et al. (2025), the use of CVAE for parameter inference of microlensed GW events was investigated, achieving rapid parameter inference in two parameters, source’s dimensionless location parameter $y$ and time delay $\Delta t_{\rm d}$, with the fastest inference speed being $5$ orders of magnitude faster than the Bayesian methods such as Bilby. In this work, we focus on microlensing events, using a point-mass lens model to describe the lensing effect on GWs, and explore the use of neural spline flows (NSFs) for parameter inference in 13 dimensions (including both sources and lenses information, i.e., lens mass $M_{\rm L}$, angular diameter distance $D_{\rm L}$, lens redshift $Z_{\rm L}$, source chirp mass $\mathcal{M}_{\rm c}$, mass ratio $q$, luminosity distance $D_{\rm L}$, source redshift $Z_{\rm S}$, merger phase $\phi_{\rm c}$, merger time $t_{\rm c}$, orbital inclination $\iota$, right ascension $ra$, and declination $dec$).
GW sources (such as the merger of binary black holes) typically exhibit spin effects Ajith et al. (2011); Callister et al. (2022); Biscoveanu et al. (2022); Fishbach et al. (2022), and studying their spins is of great significance Roulet et al. (2021); Biscoveanu et al. (2021); Tong et al. (2022). However, traditional Bayesian methods encounter significant challenges when dealing with waveforms containing spin parameters Veitch et al. (2015); Harry et al. (2016); Shi et al. (2025). Utilizing the rapidity Du et al. (2024); Wang et al. (2024a) and generalization ability Chatterjee et al. (2021); Xia et al. (2021); Wang et al. (2024b); Xu et al. (2024); Wang et al. (2024) of the NSF network can also provide a solution to this difficulty.
@@ II.1 The theory of gravitational lensing
$\xi_{0}=\sqrt{\frac{4GM_{\rm L}^{2}}{(1+Z_{\rm L})c^{2}}\cdot\frac{D_{\rm L}D_{\rm LS}}{D_{\rm S}}}.$ (1)
@@
$\Delta t_{\rm d}=2M_{\rm L}\left\{y(y^{2}+4)^{1/2}+2\ln\left[\frac{(y^{2}+4)^{1/2}+y}{(y^{2}+4)^{1/2}-y}\right]\right\},$ (4)
are the magnification factors and time delay for the two “images” respectively. Thus, the impact of the gravitational lensing effect can be determined by $y$ and $M_{\rm L}$, and further reflected by transmission factor $F(f)$.
@@ II.2 Network architecture
This work proposes a parameter inference network based on deep residual feature extraction and NSFs for efficiently reconstructing the 13-dimensional posterior distribution of physical parameters from simulated multi-detector lensed GW signals. The network is composed of a feature encoder and a probabilistic decoder cascaded together.
@@ II.3 Data sets
        $dec$ Ecliptic latitude of the binary system $\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$ $\mathrm{rad}$ Cosine
        $ra$ Ecliptic longitude of the binary system $(0,2\pi)$ $\mathrm{rad}$ Uniform
        $\psi_{\rm s}$ Polarization angle of the gravitational wave $(0,2\pi)$ $\mathrm{rad}$ Uniform
@@
        $\eta$ Location of the binary system $(10^{-12},5\times 10^{-7})$ $\mathrm{Mpc}$ Uniform
In addition, the prior of $d_{\rm L}$ from the source to observer and $D_{\rm L}$ can be obtained from $Z_{\rm L}$, and both follow a power-law distribution. The probability density function for $Z_{\rm L}$ and $Z_{\rm S}$ are as follows: $p(Z_{\rm L})\propto 3Z_{\rm L}^{2}$, and $p(Z_{\rm S}|Z_{\rm L})\propto 3/(1-Z_{\rm L}^{3})Z_{\rm S}^{2}$. To better cover the parameter space and achieve uniform sampling at different orders of magnitude, the selected $M_{\rm L}$ and $\mathcal{M}_{\rm c}$ follow a log-uniform distribution. The selection of $M_{\rm L}$ and $\eta$ range mainly refers to Ref. Kim et al. (2022). It is indicated in Ref. Takahashi and Nakamura (2003) that when redshift mass of the lens $M_{Lz}=(1+Z_{\rm L})M_{\rm L}$ and the frequency $f$ of GW satisfy the condition that $\omega\equiv 8\pi GM_{Lz}f/c^{3}\gtrsim 1$, transmission factor $F(f)$ in the wave optics (WO) asymptotically converges to the GO limit. For the cases $M_{\rm L}>10^{3}$ ${\rm M}_{\odot}$ and $f>20$ ${\rm Hz}$ considered in our study (correspondingly $\omega>7.5$), GO approximation is applicable within a certain range. Although the most rigorous approach would be to directly adopt the analytical form of $F(f)$ Takahashi and Nakamura (2003), for the purpose of exploratory verification and considering the computational cost and practicality, we chose a more concise GO approximation result. For discussions on the applicable range of the geometric optical approximation in the lensing effect of GWs, refer to Refs. Meena and Bagla (2020); Bulashenko and Ubach (2022); Bondarescu et al. (2023); Takahashi (2004).
@@ III Results and discussion
We train the NSF network using the architecture described in Section  II.2. After $316$ epochs, the loss function achieves a stable evolution, as shown in Fig. 3. The loss value of training set eventually converges to about $-11.6$, and the training loss and validation loss remain close throughout, with no sign of overfitting. At the end of training, we save the network of the epoch with the best performance, that is, the one with the lowest validation loss, and use it for testing.
We randomly select $200$ events from the test set to verify the robustness of the NSF network’s inference for selected 13 dimensions parameters. By injecting $200$ simulated lensed GW waveforms, we measure the frequency of the true parameters at a series of confidence levels. We also calculate the cumulative probability of the percentile of each parameter’s posterior distribution at the injected parameter value, and then plot the cumulative distribution function (CDF) of the joint distribution of the parameters in Fig. 4. Ideally, the injected parameter CDF should show a diagonal pattern because the CDF represents the cumulative sum of the distribution of each event. The closeness of the curve to the diagonal indicates the accuracy of the parameter inference. As shown in the figure, the curve closely follows the diagonal, and the entire curve is within the $95\%$ confidence interval, indicating that the NSF network perform excellently in accurately estimating parameters. The $p$-value obtained from the posterior distribution of the network is also marked in the legend.
@@
Figure 5 shows a comparison of the posterior inference of a lensed GW event in the spinless test set sample between our trained NSF network and Bilby dynesty, where the prior settings of Bilby dynesty are the same as those in Table  1. For the 13 parameters we studied, the NSF network successfully completed the posterior inference in less than $1$ s and provided relatively accurate posterior estimates for the lens parameters $M_{\rm L}$, $D_{\rm L}$, and $Z_{\rm L}$, which are of the greatest concern to us. Meanwhile, it also gave good inference results for some parameters of the source, such as $\mathcal{M}_{\rm c}$, $D_{\rm L}$, $Z_{\rm S}$, and $\iota$, and accurately identifying the degeneracy among these parameters. For the parameter $\iota$, NSFs provided a bimodal distribution, which was not obvious when compared with the results of Bilby dynesty in the figure. NSFs performed poorly for the sky location parameters and $\psi_{\rm s}$, $\phi_{\rm c}$, $t_{\rm c}$, and we believe that this is due to the further modulation of GWs by the lensing effect, which increases the difficulty of extracting features related to the source, preventing NSF network from extracting these parameter features well. Overall, the two-dimensional posterior distribution given by NSFs is relatively diffuse. We believe this is because the network did not extract the high-order mode features Estellés et al. (2022a, b); Mateu-Lucena et al. (2022); Diao et al. (2025) introduced by the IMRPhenomTPHM model well. We also noticed that the network did not respond well to the multimodal of $\psi_{\rm s}$, which is also an improvement direction for our network Ziegler and Rush (2019); Dinh et al. (2020); Qin et al. (2025) compared to Bilby dynesty.
In contrast, the inference accuracy of Bilby dynesty for the parameters $M_{\rm L}$, $D_{\rm L}$, $Z_{\rm L}$, $\mathcal{M}_{\rm c}$, $D_{\rm L}$, $Z_{\rm S}$, and $\iota$ is not better than that of the NSF network. Due to the introduction of higher-order modes of the waveform, the two-dimensional posterior distribution given by Bilby dynesty is indeed more concentrated. Although it had relatively better accuracy in inferring other parameters, the time it took is four orders of magnitude longer than that of NSFs. In our $10$ tests (see Table 3), the posterior inference speed of Bilby dynesty is the fastest at $60070$ s, while the inference speed of the NSF network averaged approximately $0.8$ s. This makes it possible for NSFs to apply in real-time GW signals search.
Table 3: The statistical results of the time taken for Bilby dynesty and NSFs network to independently complete the posterior inference of 13-dimensional parameters, respectively, after $10$ tests conducted.
@@
Figure 6 shows the posterior inference performance of the NSF network on the test set that includes different spin parameters. This figure depicts the distribution of error bars for $p$-values under $24$ spin values. For each spin value, we randomly select one sample from the corresponding test set (with a sample size of 2000) and perform ten random samplings (each sampling containing 200 events) to calculate ten $p$-values. These original $p$-values are represented by light gray dots in the figure. To more clearly show the trend of network performance with respect to spin values and to eliminate the influence of injected noise, we divide $p$-values under $24$ spin values into groups of $4$ for binning statistics and draw error bars. The center position of the error bars represents the mean of $p$-values in this group, and their length indicates the range of the $95\%$ confidence interval. The results show that when the spin value is less than approximately $0.32$, $p$-values are generally higher than $0.05$, indicating that the NSF network has good generalization ability within this range. However, as the spin value increases, $p$-values show an overall downward trend. When spin value is greater than $0.33$, $p$-values are significantly lower than $0.05$, which can be regarded as the generalization boundary of the NSF network for the assumed spin value scenario under the current training conditions.
Figure 6: The KS test evaluating the performance of the NSF network on the test set with added spins. We assume equal spins for the two black holes ($\chi_{\rm 1z}=\chi_{\rm 2z}$), which increase incrementally from 00 to $0.35$. The dot at the center of each error bar represents the mean $p$-value after uniformly binning the 24 datasets. The blue line illustrates the trend of these $p$-values across spin values. The light gray dots in the background represent $p$-values separately calculated for the 24 spin values, and the color shade indicates the degree of overlap of $p$-values. The black dotted line indicates a significance level corresponding to a $p$-value of $0.05$.

### v2

@@ Abstract
When gravitational waves (GWs) propagate near massive objects, they undergo gravitational lensing that imprints lens model dependent modulations on the waveform. This effect provides a powerful tool for cosmological and astrophysical studies. Due to the added parameters of lenses and the uncertainty of lens models, parameter inference for lensed GW events using traditional methods is extremely time-consuming, thus requiring more efficient parameter inference methods. In this work, we explore the use of neural spline flows (NSFs) for posterior inference of microlensed GWs, and successfully apply NSFs to the inference of 11-dimensional lens parameters. Our results demonstrate that compared with traditional methods like Bilby dynesty that rely on Bayesian inference, the NSF network we built not only achieves inference accuracy comparablye to traditional methods for the main parameters ($M_{\rm L}$, $y$, $\mathcal{M}_{\rm c}$, $d_{\rm L}$), but also can reduce the inference time from approximately 3 days to 0.8 s on average. Additionally, the network exhibits strong generalization for the spin parameters of GW sources. It is anticipated to become a powerful tool for future low-latency searches for lensed GW signals.
@@ I Introduction
Since the first direct detection of gravitational waves (GWs) in 2015, LIGO-Virgo-KAGRA (LVK) network have reported $93$ GW events in its first three observing runs, which enables us to conduct various novel tests of general relativity Abbott et al. (2016, 2018); Abbott et al. (2019a, b, 2021a) and cosmological researches The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2023). With the further analysis of the data observed in the fourth observing run of the LVK network and the construction of various high-performance gravitational wave detectors in the future (such as the Laser Interferometer Space Antenna (LISA) Barausse et al. (2020), Taiji Ruan et al. (2020), TianQin Luo et al. (2025), Einstein Telescope (ET) Maggiore et al. (2020), and Cosmic Explorer (CE) Evans et al. (2021)), the number of detected GW events will increase sharply, providing a brand-new powerful tool for research in fields such as fundamental physics, cosmology, and astrophysics. Among these GW signals, it is particularly worthy of attention those signals affected by the lensing effect. The lensing effect of GWs is caused by celestial bodies located between detectors and sources, which will result in a significant modulation of GW signals Dai and Venumadhav (2017); Dai et al. (2018).
@@
Recently, multiple studies have conducted numerous searches for the GW lensing effect, but none have clearly detected lensed GW signals. The methods employed include traditional search techniques based on matched filtering Abbott et al. (2021b); Abbott et al. (2024); Janquart et al. (2023); Barsode et al. (2025); Goyal et al. (2024) and those incorporating deep learning technology Goyal et al. (2021); Kim et al. (2022); Schäfer et al. (2023). Nevertheless, as more and more GW events are expected to be reported, and given the predictions regarding the detection rate Li et al. (2018); Ng et al. (2018); Wang et al. (2021b); Xu et al. (2022); Brando et al. (2025), we remain hopeful that GW signals affected by the gravitational lensing effect will be detectable in future observational runs.
@@
Parameter inference for lensed GW events is a crucial step in studying the GW lensing effect. The traditional parameter inference process based on Bayesian methods is extremely time-consuming, especially for lensed GW events with additional lens parameters, which significantly reduces its efficiency. Recent studies have shown that deep learning techniques such as Conditional Variational Autoencoders (CVAE) Gabbard et al. (2022) and Normalizing Flows (NFs) Green and Gair (2021); Green et al. (2020); Dax et al. (2021); Shen et al. (2022); Whittaker et al. (2022); Dax et al. (2023); Zhao et al. (2023); Sun et al. (2024a); Du et al. (2024); Sun et al. (2024b); Leyde et al. (2024); Dax et al. (2025); Xiong et al. (2025) can effectively and rapidly perform parameter inference, providing powerful tools for low-latency GW signal searches. In Ref. Nerin et al. (2025), the use of CVAE for parameter inference of microlensed GW events was investigated, achieving rapid parameter inference in two parameters, source’s dimensionless location parameter $y$ and time delay $\Delta t_{\rm d}$, with the fastest inference speed being $5$ orders of magnitude faster than the Bayesian methods such as Bilby. In this work, we focus on microlensing events, using a point-mass lens model to describe the lensing effect on GWs, and explore the use of neural spline flows (NSFs) for parameter inference in 11 dimensions (including both sources and lenses information, i.e., lens mass $M_{\rm L}$, angular diameter distance of lens $D_{\rm L}$, source chirp mass $\mathcal{M}_{\rm c}$, mass ratio $q$, luminosity distance of source $d_{\rm L}$, location parameter of the source $y$, merger phase $\phi_{\rm c}$, merger time $t_{\rm c}$, orbital inclination $\iota$, right ascension $ra$, and declination $dec$).
GW sources (such as the merger of binary black holes) typically exhibit spin effects Ajith et al. (2011); Callister et al. (2022); Biscoveanu et al. (2022); Fishbach et al. (2022), and studying their spins is of great significance Roulet et al. (2021); Biscoveanu et al. (2021); Tong et al. (2022). However, traditional Bayesian methods encounter significant challenges when dealing with waveforms containing spin parameters Veitch et al. (2015); Harry et al. (2016); Shi et al. (2025). Utilizing the rapidity Du et al. (2024); Wang et al. (2024a) and generalization ability Chatterjee et al. (2021); Xia et al. (2021); Wang et al. (2024b); Xu et al. (2024); Wang et al. (2024c) of the NSF network can also provide a solution to this difficulty.
@@ II.1 The theory of gravitational lensing
$\xi_{0}=\sqrt{\frac{4GM_{\rm L}}{c^{2}}\cdot\frac{D_{\rm L}D_{\rm{LS}}}{D_{\rm S}}}.$ (1)
@@
$\Delta t_{\rm d}=\frac{2GM_{\rm Lz}}{c^{3}}\left\{y(y^{2}+4)^{1/2}+2\ln\left[\frac{(y^{2}+4)^{1/2}+y}{(y^{2}+4)^{1/2}-y}\right]\right\},$ (4)
are the magnification factors and time delay for the two “images” respectively. $M_{\rm Lz}=M_{\rm L}(1+Z_{\rm L})$ is redshift mass of the lens. Thus, the impact of the gravitational lensing effect can be determined by $y$ and $M_{\rm L}$, and further reflected by transmission factor $F(f)$.
@@ II.2 Network architecture
This work proposes a parameter inference network based on deep residual feature extraction and NSFs for efficiently reconstructing the 11-dimensional posterior distribution of physical parameters from simulated multi-detector lensed GW signals. The network is composed of a feature encoder and a probabilistic decoder cascaded together.
@@ II.3 Data sets
$\delta$ Ecliptic latitude of the binary system $\left(-\frac{\pi}{2},\frac{\pi}{2}\right)$ $\mathrm{rad}$ Cosine
$\alpha$ Ecliptic longitude of the binary system $(0,2\pi)$ $\mathrm{rad}$ Uniform
$\psi_{\rm s}$ Polarization angle of the gravitational wave $(0,\pi)$ $\mathrm{rad}$ Uniform
@@
$\eta$ Location of the binary system $(10^{-6},0.5)$ $\mathrm{pc}$ Uniform
In addition, the prior of $d_{\rm L}$ from the source to observer and $D_{\rm L}$ can be obtained from $Z_{\rm L}$, and both follow a power-law distribution. The probability density function for $Z_{\rm L}$ and $Z_{\rm S}$ are as follows: $p(Z_{\rm L})\propto 3Z_{\rm L}^{2}$, and $p(Z_{\rm S}|Z_{\rm L})\propto 3/(1-Z_{\rm L}^{3})Z_{\rm S}^{2}$. To better cover the parameter space and achieve uniform sampling at different orders of magnitude, the selected $M_{\rm L}$ and $\mathcal{M}_{\rm c}$ follow a log-uniform distribution. The selection of $M_{\rm L}$ and $\eta$ range mainly refers to Ref. Kim et al. (2022). It is indicated in Ref. Takahashi and Nakamura (2003) that when $M_{\rm Lz}$ and the frequency $f$ of GW satisfy the condition that $\omega\equiv 8\pi GM_{\rm Lz}f/c^{3}\gtrsim 1$, transmission factor $F(f)$ in the wave optics (WO) asymptotically converges to the GO limit. For the cases $M_{\rm L}>10^{3}$ ${\rm M}_{\odot}$ and $f>20$ ${\rm Hz}$ considered in our study (correspondingly $\omega>7.5$), GO approximation is applicable within a certain range. Although the most rigorous approach would be to directly adopt the analytical form of $F(f)$ Takahashi and Nakamura (2003), for the purpose of exploratory verification and considering the computational cost and practicality, we chose a more concise GO approximation result. For discussions on the applicable range of the geometric optical approximation in the lensing effect of GWs, refer to Refs. Meena and Bagla (2020); Bulashenko and Ubach (2022); Bondarescu et al. (2023); Takahashi (2004).
@@ III Results and discussion
We train the NSF network using the architecture described in Section  II.2. After $257$ epochs, the loss function achieves a stable evolution, as shown in Fig. 3. The loss value of training set eventually converges to about $-4.7$, and the training loss and validation loss remain close throughout, with no sign of overfitting. At the end of training, we save the network of the epoch with the best performance, that is, the one with the lowest validation loss, and use it for testing.
We randomly select $200$ events from the test set to verify the robustness of the NSF network’s inference for selected 11 dimensions parameters. By injecting $200$ simulated lensed GW waveforms, we measure the frequency of the true parameters at a series of confidence levels. We also calculate the cumulative probability of the percentile of each parameter’s posterior distribution at the injected parameter value, and then plot the cumulative distribution function (CDF) of the joint distribution of the parameters in Fig. 4. Ideally, the injected parameter CDF should show a diagonal pattern because the CDF represents the cumulative sum of the distribution of each event. The closeness of the curve to the diagonal indicates the accuracy of the parameter inference. As shown in the figure, the curve closely follows the diagonal, and the entire curve is within the $95\%$ confidence interval, indicating that the NSF network perform excellently in accurately estimating parameters. The $p$-value obtained from the posterior distribution of the network is also marked in the legend.
@@
Figure 5 shows a comparison of the posterior inference of a lensed GW event in the spinless test set sample between our trained NSF network and Bilby dynesty, where the prior settings of Bilby dynesty are the same as those in Table  1. For the 11 parameters we studied, the NSF network successfully completed the posterior inference in less than $1$ s and provided relatively accurate posterior estimates for the lens parameters $M_{\rm L}$ and $y$, which are of the greatest concern to us. Here, $M_{\rm L}$ refers to the lens mass in the source frame (not the red-shifted mass). Meanwhile, it also gave good inference results for some parameters of the source, such as $\mathcal{M}_{\rm c}$ and $d_{\rm L}$. For parameters $\iota$, $q$, $\delta$ and $\alpha$, NSFs provided a bimodal distribution, which was not obvious when compared with the results of Bilby dynesty in the figure. NSFs performed poorly for $\psi_{\rm s}$, $\phi_{\rm c}$ and $t_{\rm c}$, and we believe that this is due to the further modulation of GWs by the lensing effect, which increases the difficulty of extracting features related to the source, preventing NSF network from extracting these parameter features well. Overall, the two-dimensional posterior distribution given by NSFs is relatively diffuse. We believe this is because the network did not extract the high-order mode features Estellés et al. (2022a, b); Mateu-Lucena et al. (2022); Diao et al. (2025) introduced by the IMRPhenomTPHM model well.
In contrast, the inference accuracy of Bilby dynesty for the parameters $M_{\rm L}$, $y$, $\mathcal{M_{\rm c}}$ and ${d}_{\rm L}$ is not better than that of the NSF network. Due to the introduction of higher-order modes of the waveform, the two-dimensional posterior distribution given by Bilby dynesty is indeed more concentrated. Although it had relatively better accuracy in inferring other parameters, the time it took is four orders of magnitude longer than that of NSFs. In our $10$ tests (see Table 3), the posterior inference speed of Bilby dynesty is the fastest at $60070$ s, while the inference speed of the NSF network averaged approximately $0.8$ s. This makes it possible for NSFs to apply in real-time GW signals search.
Table 3: The statistical results of the time taken for Bilby dynesty and NSFs network to independently complete the posterior inference of 11-dimensional parameters, respectively, after $10$ tests conducted.
@@
Figure 6 shows the posterior inference performance of the NSF network on the test set that includes different spin parameters. This figure depicts the distribution of error bars for $p$-values under $36$ spin values. For each spin value, we randomly select one sample from the corresponding test set (with a sample size of 2000) and perform $16$ random samplings (each sampling containing 200 events) to calculate $16$ $p$-values. These original $p$-values are represented by light gray dots in the figure. To more clearly show the trend of network performance with respect to spin values and to eliminate the influence of injected noise, we divide $p$-values under $36$ spin values into groups of $4$ for binning statistics and draw error bars. The center position of the error bars represents the mean of $p$-values in this group, and their length indicates the range of the $95\%$ confidence interval. The results show that when the spin value is less than approximately $0.17$, $p$-values are generally higher than $0.05$, indicating that the NSF network has good generalization ability within this range. However, as the spin value increases, $p$-values show an overall downward trend. When spin value is greater than $0.17$, $p$-values are significantly lower than $0.05$, which can be regarded as the generalization boundary of the NSF network for the assumed spin value scenario under the current training conditions.
Figure 6: The KS test evaluating the performance of the NSF network on the test set with added spins. We assume equal spins for the two black holes ($\chi_{\rm 1z}=\chi_{\rm 2z}$), which increase incrementally from 00 to $0.35$. The dot at the center of each error bar represents the mean $p$-value after uniformly binning the 36 datasets. The blue line illustrates the trend of these $p$-values across spin values. The light gray dots in the background represent $p$-values separately calculated for the 36 spin values, and the color shade indicates the degree of overlap of $p$-values. The black dotted line indicates a significance level corresponding to a $p$-value of $0.05$.